In a "Probability and Statistics for CS" class with 160 students enrolled in it, by the
time of the final exam, 65% of the students have mastered discrete probabilities
70% have mastered continuous probabilities, and 45% have mastered both. For a
student to pass the class with at least a C, the student must be able to use at least
one type of probabilities. How many students passed the class?
Answer #1

P(Disc Prob) = 0.65

P(Con Prob) = 0.70

P(Disc & Con Prob) = 0.45

P(Disc Prob U Con Prob) = P(Disc Prob) + P(Con Prob) - P(Disc & Con Prob)

P(Disc Prob U Con Prob) = 0.65 + 0.70 - 0.45

P(Disc Prob U Con Prob) = 0.90

No of students = 160 * 0.90 = 144

ANS : 144 students passed the class